# Learning Manipulation Skills through Robot Chain-of-Thought with Sparse Failure Guidance

### v2

@@ Abstract
Defining reward functions for skill learning has been a long-standing challenge in robotics. Recently, vision-language models (VLMs) have shown promise in defining reward signals for teaching robots manipulation skills. However, existing works often provide reward guidance that is too coarse, leading to inefficient learning processes. In this paper, we address this issue by implementing more fine-grained reward guidance. We decompose tasks into simpler sub-tasks, using this decomposition to offer more informative reward guidance with VLMs. We also propose a VLM-based self imitation learning process to speed up learning. Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP. Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rate compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
Keywords: Manipulation Skills, Vision Language Models, Chain-of-Thought
1 Introduction
@@
Defining reward functions for learning robotic skills has been a long standing challenge in robotics [1]. Traditional methods typically define the rewards through human knowledge and require significant engineering efforts. Some methods also use expert demonstrations to define rewards [1]. However, this procedure requires expert supervision and can be infeasible in complex scenarios [2] . Recently, the success of vision language models (VLM) has inspired researchers to leverage its power to define rewards for skill learning. For example, RoboCLIP [2] captures the temporal information from robot movements with a video backbone S3D [3].
@@
In spite of the preliminary success of VLM-based reward modeling, a downside of these existing works is that their language guidance is too coarse for learning. Consider the task of opening the door, existing works only use the language embedding of "opening the door" in reward definition to guide the learning process. Such guidance can be too sparse for learning: when the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened. In contrast, in human world we typically use more fine-grained guidance such as "move towards the door", "move the hand towards the handle" and "grab the handle" during the teaching process. Intuitively, using such decomposition can make the guidance more precise, and will lead to richer and denser feedback.
@@
Based on this intuition, in this paper we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem. Our main insight is to decompose the task description into a series of actionable steps and then leverage VLM to define a fine-grained reward to guide skill learning, which can make the learning process more sample efficient and yield higher success rates. We also propose a VLM-based self-imitation procedure to further speed up the learning process. In the experiments, we demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline. Ablation studies show that each component of our algorithm contributes to the overall performance.
@@
Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using a robot chain-of-thought processing, breaking it down into several detailed prompts. The robot also considers some failure guidance. Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills. Successful experiences are recorded by foundation models to reinforce effective behaviors (i.e., VLM-based self-imitation).
2 Related Work
2.1 Learning Manipulation Skills for Robotics
Acquiring general manipulation abilities can be facilitated through learning from expert demonstrations, using behavioral cloning (BC) [4, 5, 6] or inverse reinforcement learning (IRL) [7, 8]. BC directly maps the state space to the action space to mimic the expert’s actions. However, BC faces several limitations, including compounding errors [9], the copycat problem [10], and out-of-distribution sample shifts [11]. In contrast, IRL derives both the reward function and the policy from expert demonstrations. However, IRL struggles with sample efficiency, particularly in real-world applications [12, 13, 14].
Mastering manipulation skills with language inputs directly presents several challenges, such as a lack of reward signals and inefficiency in understanding vision-language contexts. Recent studies have explored integrating VLMs to enhance the understanding of language instructions [15]. Our algorithm is designed by effectively leveraging VLMs as a source of rewards.
2.2 Foundation Models for Robotics
@@
VLMs equip robot agents with compact pre-trained representations. R3M [16] co-trains video and language instructions in a self-supervised manner [17], enhancing the robot’s ability to mimic expert actions. PaLM-E [18] integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process. RT-1 [19] uses transformers to better mimic expert behaviors. RoboFlamingo [20] uses multi-step observation inputs to capture a comprehensive view of observational data. RT-2 [21] innovates by conceptualizing robot actions as a form of language. Furthermore, RT-H [22] introduces an action hierarchy through language motions, leveraging the chain-of-thought reasoning capabilities of large language models.
@@
VLMs significantly enhance robot manipulation capabilities by offering valuable reward signals. CLIP [23], for instance, aligns image representations with language embeddings. Consequently, [24] employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal. LIV [25] innovatively combines VIP [26] and CLIP to more effectively learn both visual and linguistic goal representations. RoboCLIP [2] adopts the S3D [3] representation model to assess the reward for the final observation of a trajectory. Furthermore, RL-VLM-F [27] directly utilizes GPT-4V to label preferences for agent rollout data.
However, current models still struggle to master a broad spectrum of robot tasks, often succeeding only in simple tasks such as closing a drawer. VLMs like CLIP [23] and LIV [25] lack an understanding of the manipulation process. On the other hand, RoboCLIP [2] lacks spatial reasoning capabilities and task completion awareness. RL-VLM-F [27] excels in spatial reasoning for manipulation tasks but is deficient in process understanding capabilities. In contrast, our algorithm not only comprehends detailed language instructions from the robot chain-of-thought but also considers failure scenarios. Additionally, it leverages VLM’s task completion reasoning ability through self imitation process, offering a more comprehensive approach to master robot tasks.
3 Background
@@
Chain-of-Thought (CoT) The CoT [28] approach is a reasoning method designed to enhance the problem-solving capabilities of large language models. This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving. By explicitly detailing these intermediate steps, large language models such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
@@ 4 Method
Figure 2: First, text embeddings are generated using robot chain-of-thought (positive prompts) and human failure guidance (negative prompts). Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings. Finally, VLMs evaluate the final observation. If this observation is deemed successful, this trajectory is recorded in the success replay buffer for self imitation.
@@
In section 4.1, we introduce the robot chain-of-thought interpretation. In section 4.2, we explain how to define the rewards for policy learning with VLMs. Finally, in section 4.3, we introduce VLM-based self imitation procedure via success experience regularization.
4.1 Robot Chain-of-Thought
@@
Manipulation skills consist of a series of actionable control steps. For instance, opening a door includes sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door. Decomposing tasks simplifies the manipulation process, affording robots a greater level of understanding and control.
@@
We use GPT-4V to interpret task instructions such as "open the door." With GPT-4V, this instruction is broken down into a series of sub-goals $l_{i}^{+}$ through a CoT process, such as "the robot hand approaches the door’s handle," "the robot hand grasps the door’s handle," "the robot hand pulls the door’s handle back," and "the robot hand opens the door." Consequently, we leverage the fine-grained task descriptions from GPT-4V interpretation to enhance the agent’s understanding of the tasks.
4.2 VLMs as Rewarder
A central challenge we address is defining a reward function that effectively reinforces desired behaviors. In this work, we argue that VLMs can provide the agent with a much more fine-grained reward guidance. Consequently, we propose a novel reward design mechanism based on robot CoT. Specifically, we explore the usage of VideoCLIP [29] as a VLM. VideoCLIP excels in capturing the temporal dynamics of video segments and interpreting associated language prompts, aiming to align robot movements with actions described in the task descriptions.
The reward signal for learning manipulation skills is defined as follows:
$\text{reward}=\frac{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
@@ 4.2 VLMs as Rewarder
where $z_{v}$ is the video segment embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot CoT interpretation $l^{+}$, and $z_{l}^{-}$ is a set of potential failure case description embeddings guided by humans. Here, offline human failure guidance helps the agent consider potential failure cases, improving the policy for robustly learning manipulation skills. To enable the agent to recognize and adjust to potential setbacks, we provide sparse human failure guidance before training with 2 to 4 prompts. The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims to distinguish between video segments and corresponding positive or negative texts. We argue that a similar NCE-based reward formulation can leverage the VideoCLIP model rather than the reward definition in RoboCLIP [2]. Moreover, Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent. Otherwise, the reward signal would penalize undesired behaviors.
4.3 VLM-based Self Imitation
@@
Our reward design mechanism motivates the agent to explore the necessary manipulation skills. Empirically, we find that the VideoCLIP rewards don’t reflect the completion of the manipulation tasks well. They focus mostly on whether the robot’s movements align with detailed language goals. To better reinforce task completion and speed up the learning, we introduce VLM-based self imitation procedure.
@@
We employ GPT-4V to assess the success of the agent’s final observation in its rollout trajectory. If it deemed successful, this trajectory is archived into a success replay buffer and we correct the reward function as: $r(s_{T})=r(s_{T})+100$. Here, $s_{T}$ is the successful final observation of the rollout trajectory. Subsequently, we introduce a regularization mechanism aimed at refining the agent’s policy to reinforce successful behavior. This motivates the agent to reinforce the success experience. The regularization loss is integrated with the policy loss, formulated as follows:
@@
where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter. By incorporating success experiences into the policy loss calculation, the agent is encouraged to repeat desired behaviors.
5 Experiment
@@
In this section, we assess the effectiveness of our proposed algorithm RoboCoT. The experiments are designed to address two questions: (1) How does our algorithm perform compared to other state-of-the-art algorithms? The results are presented in section 5.1. (2) What are the contributing factors in our algorithm? The ablation studies are presented in section 5.2.
@@
Environments To evaluate the effectiveness of our proposed algorithm RoboCoT, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment. We select 10 tasks from simple to complex, including drawer, window, faucet, button press, door, and soccer, as described by [30]. Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
@@
Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 [31] as the reinforcement learning algorithm. Detailed implementation specifics can be found in the references: CLIP [24], LIV [25], and RoboCLIP [2]. All experiments, including those involving our proposed algorithm, receive the same language instructions as inputs. Implementation details are provided in the Appendix 8.
5.1 Overall Comparison
Figure 3: Training curves for baselines and our algorithm within 3 random seeds. Each data point is evaluated with 20 sampled trajectories. The shaded area displays the range of one standard deviation. Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × improvement compared to RoboCLIP.
To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing it with established baselines: CLIP [24], LIV [25], and RoboCLIP [2]. CLIP [24] employs CLIP [23] as the vision-language model to measure the distance between current observations and language goals. LIV [25] aligns goal representations with language instructions, and RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward. However, CLIP [23] and LIV [25] lack an fine-grained reward guidance, while RoboCLIP [2] lacks task completion awareness. Figure 3 illustrates the training curves, depicting the success rates for each algorithm. Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.

### v3

@@ Abstract
Defining reward functions for skill learning has been a long-standing challenge in robotics. Recently, vision-language models (VLMs) have shown promise in defining reward signals for teaching robots manipulation skills. However, existing work often provides reward guidance that is too coarse, leading to insufficient learning processes. In this paper, we address this issue by implementing more fine-grained reward guidance. We decompose tasks into simpler sub-tasks, using this decomposition to offer more informative reward guidance with VLMs. We also propose a VLM-based self imitation learning process to speed up learning. Empirical evidence demonstrates that our algorithm consistently outperforms baselines such as CLIP, LIV, and RoboCLIP. Specifically, our algorithm achieves a 5.4×5.4$\times$5.4 × higher average success rates compared to the best baseline, RoboCLIP, across a series of manipulation tasks.
I INTRODUCTION
@@
Defining reward functions for learning robotic skills has been a long-standing challenge in robotics [1]. Traditional methods typically define rewards through human knowledge and require significant engineering efforts. Some methods also use expert demonstrations to define rewards [1]. However, this procedure requires expert supervision and can be infeasible in complex scenarios [2] . Recently, the success of vision language models (VLM) has inspired researchers to use its power to define rewards for skill learning. For example, RoboCLIP [2] captures the temporal information from robot movements with an S3D video backbone [3], therefore defining the rewards for skill learning.
@@
In spite of the preliminary success of VLM-based reward modeling, a downside of these existing works is that their language guidance is too coarse for learning. Consider the task of opening the door, existing works only use the language embedding of "opening the door" in reward definition to guide the learning process. Such guidance can be too coarse for learning: When the robot is not able to approach the door at the beginning of learning process, the rewarder will always provide a near-zero feedback since the door is not likely to be opened. In contrast, in human world we typically use more fine-grained guidance such as "move towards the door", "move the hand towards the handle" and "grab the handle" during the teaching process. Intuitively, using such decomposition [4, 5, 6, 7, 8] can make the guidance more precise, and will lead to richer and denser feedback.
@@
Based on this intuition, in this paper, we present a novel framework named Rewarder from Robot Chain-of-Thought with Sparse Failure Guidance (RoboCoT) to address this problem. Our main insight is to decompose the task description into a series of actionable steps and then leverage VLM to define a fine-grained reward to guide skill learning, which can make the learning process more sample efficient and yield higher success rates. In addition, we propose a VLM-based self-imitation procedure to further speed up the learning process. In the experiments, we demonstrate that our approach performs well across a broad spectrum of tasks, achieving a 5.4×5.4$\times$5.4 × performance improvement over the best existing baseline. Empirical studies show that each component of our algorithm contributes to the overall performance.
@@
Figure 1: First, the robot receives the language instruction, e.g., "open the door." Subsequently, our algorithm interprets this instruction using a robot chain-of-thought processing, breaking it down into several detailed prompts. Then, the robot initiates reinforcement learning with our designed reward signal to learn the skills with the consideration about the failure guidance. Successful experiences are recorded by foundation models to reinforce effective behaviors, i.e., VLM-based self-imitation.
II RELATED WORK
Defining reward functions for robotic skill learning has long been a significant challenge. This challenge can be addressed through learning from demonstrations using inverse reinforcement learning (IRL) [9, 10, 11, 12, 13]. However, IRL requires expert supervision, which can be infeasible in complex scenarios [2]. Our RoboCoT overcomes this limitation by effectively leveraging VLMs as a source of rewards. By fully utilizing prior knowledge from VLMs, RoboCoT comprehends the manipulation process and subsequently achieves superior manipulation performance by learning directly from language instructions.
II-A Robot Chain-of-Thought for Robotics
Robot chain-of-thought [4] is employed to interpret task descriptions and enhance high-level planning performance in robotics. For example, ViLa [7] utilizes GPT-4V to decompose tasks into sub-goals based on robot observations and task descriptions, demonstrating significant success in understanding complex robotic tasks. HiP [5] achieves hierarchical planning by diffusing the future observations, therefore better decomposing the complex tasks. However, current approaches primarily apply robot chain-of-thought to high-level planning, thereby limiting its potential for comprehensively understanding robotic tasks. Our RoboCoT extends the application of robot chain-of-thought by designing more fine-grained reward signals. Empirical studies demonstrate its promise in directly following language instructions.
II-B Vision Language Models for Robotics
@@
VLMs equip robot agents with compact, pre-trained representations. R3M [14] co-trains video and language instructions in a self-supervised manner [15], enhancing the robot’s ability to mimic expert actions. PaLM-E [16] integrates visual observations with large language models, facilitating low-level robot control that includes reasoning about the control process. RT-1 [17] uses transformers to better mimic expert behaviors. RoboFlamingo [18] uses multi-step observation inputs to capture a comprehensive view of observational data. RT-2 [19] innovates by conceptualizing robot actions as a form of language. RT-H [20] introduces an action hierarchy through language motions, leveraging the robot chain-of-thought reasoning capabilities of large language models (LLMs).
@@
VLMs significantly enhance robot manipulation capabilities by offering valuable reward signals. CLIP [21], for instance, aligns image representations with language embeddings. Consequently, [22] employs CLIP to measure the reward distance between a robot’s current observation and its linguistic goal. LIV [23] innovatively combines VIP [24] and CLIP to more effectively learn both visual and linguistic goal representations. RoboCLIP [2] adopts the S3D [3] representation model to assess the reward for the final observation of a trajectory. Additionally, [25] considers potential failure cases for designing reward signals in imitating language instructions. RL-VLM-F [26] directly uses GPT-4V to label preferences for agent rollout data.
However, current models use coarse instructions like "open the door" for designing reward signals. Our RoboCoT improves upon this by leveraging the robot chain-of-thought technique to design more fine-grained reward signals for following language instructions. Empirical studies demonstrate the promise of using VLMs as rewarders in imitation learning.
III BACKGROUND
@@
Chain-of-Thought (CoT) The CoT [4] approach is a reasoning method designed to enhance the problem-solving capabilities of LLMs. This method involves decomposing a problem into a series of intermediate steps or thoughts that collectively guide towards a solution, mirroring the human approach to complex problem solving. By explicitly detailing these intermediate steps, LLMs such as GPT-4V are better equipped to navigate and resolve complex challenges that require advanced understanding and logical deduction.
@@ IV METHOD
Figure 2: First, text embeddings are generated with robot chain-of-thought (positive prompts) and sparse failure guidance (negative prompts). Next, a moving window slides across the agent’s rollout trajectory, calculating the NCE reward based on both video and text embeddings. Finally, VLMs evaluate the final observation. If this observation is deemed successful, this trajectory is recorded in the replay buffer for self imitation procedure.
@@
In Section IV-A, we present the robot chain-of-thought process for comprehending the robotic tasks. In Section IV-B, we outline the process of defining rewards for skill learning using Vision-Language Models (VLMs). Finally, in Section IV-C, we introduce the VLM-based self-imitation procedure through success experience regularization.
IV-A Robot Chain-of-Thought
@@
Robotic manipulation skills consist of a series of actionable control steps. For example, opening a door involves sub-tasks such as approaching the handle, grasping it, pulling it back, and opening the door. Decomposing tasks with a robot chain-of-thought (CoT) process simplifies the manipulation process, providing robots with a deeper understanding and greater control.
However, previous works [2, 25, 22, 23] have primarily used coarse language instructions, such as ’open the door,’ in defining reward signals. This level of guidance is insufficient for learning robust manipulation skills. Our RoboCoT addresses this limitation by employing robot chain-of-thought process to decompose tasks for more fine-grained reward design.
@@
We use GPT-4V to interpret task instructions like ’open the door.’ Through robot chain-of-thought process, GPT-4V breaks down this instruction into a series of sub-goals, such as ’the robot hand approaches the door’s handle,’ ’the robot hand grasps the door’s handle,’ ’the robot hand pulls the door’s handle back,’ and ’the robot hand opens the door.’ Consequently, we leverage these task descriptions from the GPT-4V interpretation to define a fine-grained reward signal in skill learning, thereby enhancing the agent’s understanding of the robotic tasks.
IV-B VLMs as Rewarder
Defining reward functions for robot skill learning is a long standing challenge. In this work, we leverage VLMs to design more fine-grained reward signals in skill learning. Moreover, we also consider the potential failure guidance in teaching robot manipulation skills which has been investigated in [25].
We explore the usage of VideoCLIP [27] as a rewarder. VideoCLIP excels in capturing the temporal dynamics of video slices and interpreting associated language prompts, aiming at aligning robot movements with actions described in the task descriptions.
We define the reward signal for skill learning as:
$r=\frac{\max\limits_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)}{\max\limits_{z_{l}\in\{z_{l}^{+}\}}\exp(z_{v}\cdot z_{l}/\tau)+\sum\limits_{z^{\prime}_{l}\in{\{z_{l}^{-}\}}}\exp(z_{v}\cdot z^{\prime}_{l}/\tau)},$ (3)
where $r$ represents the rewards for moving window stride timesteps, $z_{v}$ is the video embedding obtained using a moving window sliding on the agent rollout trajectory, $z_{l}^{+}$ is a set of detailed task description embeddings from robot chain-of-thought process, and $z_{l}^{-}$ is a set of potential failure case embeddings guided by humans. Here, sparse failure guidance motivates the agent to consider potential failure cases, improving the robustness.
@@ IV-B VLMs as Rewarder
The training of VideoCLIP, employing a methodology akin to the InfoNCE loss, aims at distinguishing between video slices and corresponding positive or negative texts. We find that this similar NCE-based reward formulation can better leverage the VideoCLIP model which is different from the reward definition in RoboCLIP [2]. Moreover, Equation 3 ensures that robot behaviors aligning with the detailed task descriptions and avoiding failure scenarios yield higher rewards to the agent.
Figure 3: Training curves for baselines and our algorithm within 3 random seeds. Each data point is evaluated with 20 sampled trajectories. The shaded area displays the range of one standard deviation. Our algorithm gets best performance compared to other baselines, achieves 5.4×5.4$\times$5.4 × average improvement compared to RoboCLIP.
IV-C VLM-based Self Imitation
@@
Our reward design mechanism encourages the agent to explore the necessary skills. We find that the VideoCLIP rewards focus on whether the robot movements align with the detailed language instructions. Consequently, speeding up the learning is necessary. To better speed up the learning, we introduce VLM-based self imitation procedure.
@@
We employ VLMs, i.e., GPT-4V, to assess the success of the agent’s final observation in its rollout trajectory. If it deemed successful, this trajectory is recorded into a success replay buffer $D_{E}$. Additionally, we correct the reward function on this successful observation as: $r(o_{T})=r(o_{T})+100.0$. Here, $r(o_{T})$ is the reward function for this successful final observation $o_{T}$ in the rollout trajectory. Subsequently, we introduce a regularization term in loss function, aimed at reinforcing success behaviors. The regularization loss is integrated with the policy RL loss, formulated as follows:
@@
where $D_{E}$ is the success experience replay buffer and $\lambda$ is the hyper-parameter. By incorporating success experiences into the policy loss calculation, we speed up the skill learning process. Here we note that in our experiments, $\lambda$ is set to $1$.
V EXPERIMENT
@@
In this section, we assess the effectiveness of our proposed algorithm RoboCoT. The experiments are designed to address two questions: (1) How does our algorithm RoboCoT perform compared to other state-of-the-art algorithms? The results are presented in section V-A. (2) What are the contributing factors in our RoboCoT? The ablation studies are presented in section V-B.
@@
Environments To evaluate the effectiveness of our proposed algorithm RoboCoT, we conduct experiments on a wide array of manipulation tasks using MetaWorld-v2 as the test robotic environment. We select 10 tasks from simple to complex, including drawer, window, faucet, button, door, and soccer, as described in [28]. Among these tasks, opening objects such as drawers, windows, faucets, and doors is significantly more challenging than closing them.
@@
Setup To ensure a fair evaluation across all experiments, we use DrQ-v2 [29] as the reinforcement learning algorithm. Detailed implementation specifics can be found in the references: CLIP [22], LIV [23], and RoboCLIP [2]. All the experiments, including those involving our proposed algorithm RoboCoT, receive the same language instructions as inputs.
V-A Overall Comparison
To demonstrate the superiority of our algorithm, we conduct a series of experiments comparing our RoboCoT with existing baselines: CLIP [22], LIV [23], and RoboCLIP [2].
CLIP [22] employs CLIP [21] as the VLMs to measure the distance between current observations and language goals. LIV [23] aligns VIP [24] goal representations with language instructions, therefore measuring the reward distance between current robot observation with language goals. RoboCLIP [2] utilizes S3D [3] to measure the distance between trajectory videos and language as a form of reward. Since the reward is defined with a coarse language instruction, CLIP [21], LIV [23] and RoboCLIP [2] lack fine-grained reward guidance. On the other hand, our RoboCoT uses robot chain-of-thought and sparse failure guidance to define the rewards for skill learning. Consequently, RoboCoT can better acquire the skills through reinforcement learning.
Figure 3 illustrates the training curves, depicting the success rates for each algorithm. Each evaluation involves 20 trajectories, and the experiments were replicated across three different environmental random seeds to introduce variability in the position of objects.
Figure 4: Final policy performance for ablations and our RoboCoT. The results show that each component of our RoboCoT contributes to the final performance.